OpenStudy (anonymous):

Mrs. Kiehn bought 80 feet of fence to put up in her backyard. Her backyard has four sides, all of which are equal in length. She plans on making the sides of the fence perpendicular to each other. What is the area of her yard?

ganeshie8 (ganeshie8):

hint : perimeter of square = 4*s = 80

ganeshie8 (ganeshie8):

solve above - you would get the side "s"

OpenStudy (anonymous):

400 feet square

ganeshie8 (ganeshie8):

to find area use the formula area = s^2